A particle rotates in a circle with a centripetal acceleration of 9.81 meters per second squared. If the radius is doubled without changing the particle's speed, what is the new centripetal acceleration in terms of the original value?
Answer #1

centripetal acceleration

a = v^2/r

if  radius is doubled i.e r= 2r

then

If the radius doubles, that means the bottom line of the fraction doubles, and therefore the overall value of the fraction becomes halved.

Therefore the centripetal acceleration becomes halved.

so new a = 9.81/2 = 4.905 m/s^2